# A Methodological Report on Anomaly Detection on Dynamic Knowledge Graphs

### v1

@@ 4.2 Performance of Models
The experimental results of each representation are shown in tables 1 and 2. Table 1 focuses on the performance of individual models, while table 2 highlights the performance of different ensemble methods. D1, D2, and D3 represent sequential data representation, one-hop graph representation, and two-hop graph representation, respectively.
Table 1 shows how each model performs under different structural conditions of the dataset. The results show that the inclusion of graph structures and service interactions generally improves model performance, particularly in terms of recall and F1 score.
Table 1: Performance metrics for each model across different phases. The best performances are highlighted in bold.
@@
In order to effectively combine multiple models, the second part of the experiments evaluates the performance of different ensemble methods, as shown in Table 2. Different ensemble strategies are tested, such as combining XGB, SVM, stacked layers of the Self Attention model, and Isolation Forest. In the evaluation of the ensemble methods, we have compared two different methods, soft voting, hard&unanimous voting and hard&majority voting. Table 2 shows the best performances obtained. Overall, the soft voting approach provided better results for our objectives.
Table 2: Performance metrics for ensemble methods across different phases.The best performance are highlighted in bold.
@@
XGB + SVM + SA soft D2+D3 0.51429 0.5 0.52941
The individual models using the three representations mentioned in Section 3 and the results using ensemble learning both significantly outperforms the baseline result  [^3] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
@@ 5 Conclusion
In this paper, we investigated the effectiveness of different machine learning and deep learning models on the challenging task of anomaly detection on dynamic knowledge graphs, with a specific focus on a microservices environment. Using a structured experimental approach, we analysed the performance of the models in three different representations: sequential data modelling, one-hop graph structure, and two-hop graph structure.
Our results show that incorporating graph structures significantly improves the performance of anomaly detection models. In particular, the Two-Hop Graph Structure, which includes service interactions, produced the best results across most metrics, highlighting the importance of capturing complex interdependencies within the data. Deep learning models, particularly the Two-Stacked Self-Attention (2StackedSA) model, showed superior performance in leveraging these structures.
Furthermore, the ensemble methods provided a substantial boost in performance by combining the strengths of top-performing individual models. Among the ensemble strategies, the Unanimous Voting approach, which requires all models to agree on the prediction of an anomaly, outperformed the Majority Voting approach in terms of precision. This aligns well with the competition’s emphasis on precision, highlighting the effectiveness of this conservative yet accurate method.
Future work should address several areas for improvement. Firstly, improving the scalability of our models to handle larger and more complex datasets would be crucial. Secondly, exploring more advanced ensemble techniques, such as stacking or blending, could further improve detection accuracy. Thirdly, the integration of real-time anomaly detection capabilities would be highly beneficial for practical applications in dynamic environments. Finally, our study was limited to binary classification (normal/abnormal). Future research should focus on improving the accuracy of predicting different types of anomalies. Furthermore, although we used graph structure information, we did not apply Graph Neural Networks (GNNs), which could be a promising direction for future investigation and testing.
This study demonstrates the significant potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs. By refining these methods and exploring the future directions outlined, we can develop more robust, accurate and scalable solutions for complex anomaly detection tasks in various domains.

### v2

@@ 4.2 Performance of Models
This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, one-hop, and two-hop graph representations. Table 2 summarises the performance of individual models, while table 3 focuses on the results of different ensemble methods.
The first part of the experiments evaluates the performance of each model under different data representations, as shown in Table 2. The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 score.
Table 2: Performance of each model across different data representations. The best performance is highlighted in bold.
@@
The second part of the experiments evaluates the performance of different ensemble methods, as shown in table 3. We tested combinations of XGB, SVM, Self-Attention, and Isolation Forest, along with different voting mechanisms: soft, hard-unanimous, and hard-majority. Table 3 shows that the best performance is achieved with the combination of SVM, XGB and SA in hard-unanimous voting.
Table 3: Performance of different ensemble methods.The best performance is highlighted in bold.
@@
XGB + SVM + SA hard,unanimous D2+D3 0.51429 0.5 0.52941
Figure 4: The figure shows the F1 scores of the models across the three dataset levels, as well as the F1 scores obtained after applying ensemble learning methods. The individual models include XGB, SVM, TCN, LSTM, GRU, and the Self-Attention Model. The models used in the ensemble learning approach are XGB, Isolation Forest, SVM, and the Self-Attention Model.
Fig. 4 shows that the one-hop representation, two-hop representation as well as the ensemble learning methods all significantly outperforms the baseline result  [^5] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
@@ 5 Conclusion
In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a microservices environment. Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, one-hop, and two-hop graph structures.
Our results show that incorporating graph structures significantly enhances model performance. The two-hop graph structure, which includes service interactions, delivered the best results across most metrics, emphasising the importance of capturing complex interdependencies. Deep learning models, especially the Self-Attention model, performed particularly well in leveraging these structures.
Furthermore, the ensemble methods significantly boosted performance by combining the strengths of the top individual models. Unanimous voting, which requires all models to agree on an anomaly, outperformed majority voting in terms of precision, matching the competition’s emphasis on precision.
Future work should focus on several improvements. First, improving model scalability to handle larger, more complex datasets is essential. Second, exploring advanced ensemble techniques such as stacking or blending could improve detection accuracy. Third, integrating real-time anomaly detection would benefit practical applications in dynamic environments. Additionally, our study was limited to binary classification; future research should improve the accuracy of detecting different types of anomalies. Finally, although we used graph structure information, the application of graph neural networks remains a promising direction for further investigation.
This study highlights the potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs. Refining these methods and exploring the suggested future directions can lead to more robust, accurate, and scalable solutions for complex anomaly detection in various domains.